Using weiestrass M-test, show that the following series converges uniformly.



∑n^3X^n,X belongs to[-1/3,1/3]


n=1

1
Expert's answer
2022-03-30T12:51:03-0400

ANSWER

Denote "f_{n}(x) =n^{3}x^{n}." If "\ x\in\left [ -\frac{1}{3} ,\frac{1}{3}\right ]" , then

"\left |f _{n}(x) \right |\leq\frac{n^{3}}{3^{n}}" .

Let "a_{n}=\frac{n^{3}} {3^{n}}" .

"\frac{a_{n+1}}{a_{n}}=\frac{(n+1) ^{3}\cdot3^{n}}{3^{n+1}\cdot n^{3}}= \frac{1}{3}\cdot \left ( 1+\frac{1}{n} \right )^{3}" .

"\lim _{n\rightarrow\infty}\frac{1}{3}\cdot \left ( 1+\frac{1}{n} \right )^{3}=" "\frac{1}{3}\cdot" "\lim _{n\rightarrow\infty}\ \left ( 1+\frac{1}{n} \right )^{3}=" "\frac{1}{3 }\cdot \left ( \lim_{n\rightarrow\infty}\left ( 1+\frac{1}{n} \right ) \right )^{3}=\frac{1}{3}<1"

Because   "\lim_{n\rightarrow\infty}\frac{a_{n+1}}{a_{n}}=\frac{1}{3}<1\" , then the series "\sum_{n=1}^{\infty}a_{n}" converges (by the Ratio Test).

Thus, using the Weierstrass M-Test , we make sure that the series "\sum_{n=1}^{\infty}n^{3}x^{n}" converges uniformly on "\left [ -\frac{1}{3} ,\frac{1}{3}\right ]"